Compute trend percentages for Flash Corporation total revenue and net income for the following five-year period, using year 0 as the base year. Round to the nearest full percent.

(in thousands)

Year 4

Year 3

Year 2

Year 1

Year 0

Total revenue

1629

1316

$1214

$1101

$1121

Net income

129

109

97

84

87
Answer #1
Year Total revenue Trend Percentage of total revenue Net income Trend percentage of net income
0 1,121 100% 87 100%
1 1,101 1,101/1,121 = 98% 84 84/87 = 97%
2 1,214 1,214/1,121 = 108% 97 97/87 = 111%
3 1,316 1,316/1,121 = 117% 109 109/87 = 125%
4 1,629 1,629/1,121 = 145% 129 129/87 = 148%

kindly give a positive rating if you are satisfied with the solution. do comment if you have any query, Thanks.